OpenStudy (anonymous):

A brand name has a 60​% recognition rate. Assume the owner of the brand wants to verify that rate by beginning with a small sample of 4 randomly selected consumers. What is the probability that exactly 3 of the selected consumers recognize the brand​ name?

OpenStudy (kropot72):

You can use the binomial distribution to solve this.

OpenStudy (anonymous):

\[P(x)=\frac{ n! }{ (n-x)!x! }*p ^{x}*q ^{n-x}\] That's the formula I have in my notes but I also have a chart that I don't understand how I got the numbers.

OpenStudy (kropot72):

That is the correct formula. Putting in the numbers we get: \[\large \frac{4!}{(4-3)!3!}\times0.6^{3}\times(1-0.6)^{(4-3)}=?\]

OpenStudy (anonymous):

.3456?

OpenStudy (anonymous):

I think what I'm really confused by is the binomial probabilities table they want me to make. They provided an example and I'm not comprehending how they got their numbers.

OpenStudy (kropot72):

If the table relates to this question, I assume that the table would show the probabilities of exactly 0, 1, 2, 3 and 4 of the selected consumers recognize the brand​ name.

OpenStudy (kropot72):

Your calculation of P(exactly 3 consumers) is correct.

OpenStudy (anonymous):

Hm. Okay. so P(3) on the chart would be the same as my answer above? Or am I still confused?

OpenStudy (kropot72):

Without having seen the chart, it appears that you are on the right track.

OpenStudy (anonymous):

Thanks, Kropot72! I ended up getting it correct.

OpenStudy (kropot72):

Good work! You're welcome :)